OpenStudy (anonymous):

If a+b=140 and ar+br=35. What is the value of r?

OpenStudy (anonymous):

First this you need to do is solve for r ar+br=35 r(a+b)=35 r=35/(a+b)

OpenStudy (anonymous):

The next thing is use substitution so you only have one variable a+b=140 a=140-b

OpenStudy (anonymous):

I didn't think of turning r into terms of ab. Thanks!

OpenStudy (anonymous):

Then substitute 140-b in the equation you solved for r r=35/((140-b)+b) r=35/140